OpenStudy (anonymous):

Blaise borrowed $500 for a color tv set. The APR is 18.75% find his monthly payments if he pays off the loan in 24 months

OpenStudy (anonymous):

So this is simple interest where you're dealing with I = Prt? Since you're not here, I'll just assume that. I = interest in dollars P = price of item r = rate (percentage represented as decimal) t = time (in years) I = (500)(0.1875)(2) (24 months = 2 years) I = $187.50

OpenStudy (anonymous):

$500 + $187.50 = 687.50 687.50/24 months = $28.65/month

OpenStudy (anonymous):

Monthly pymt = $25.14